Fermi National Laboratory
 

Accelerator Update for March 12, 2007 - March 14, 2007

 

Monday March 12

The day shift began with Operations monitoring the Tevatron store 5274 (45.43E30) and the Recycler stash (191.0E10), with the Antiproton Source (Pbar) not stacking because it was in access, with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Operations established beam to the Meson MTest experiment at 8:40 AM.

Operations terminated store 5274 at 9:01 AM. Operators began preparing the TeV for access.

Operations halted beam to Meson and made Meson enclosures M01 – M05 ready for access.

Operations prepared CDF and D0 for access at 9:16 AM.

MI experts began a MI-8 collimator study at 10:28 AM.

Cryo system experts reported at 10:44 AM that they had completed their repair on the Pbar dry engine flywheel.

Operators returned the MI-20 pond pump to low speed at 10:56 AM. This was turned to high speed due to earlier high temperature problems, which were fixed by FESS.

Pbar experts reported at 11:04 AM that they had completed their stacktail equalizer work.

At 11:21 AM, a Linac RF station (LRF4) tripped off. Experts reset the station.

Pbar and D0 came out of access at 11:30 AM.

At 11:34 AM, Operations reported that the Linac beam was intermittent due to problems with LRF4. The station was also causing beam position problems in Booster. An expert made adjustments to LRF4 to correct these problems. They reported that the station’s power amplifier tube was getting weak.

MI-8 collimator studies ended at 11:59 AM.

Pbar resumed stacking at a low stack rate at 12:02 PM.

Pbar experts began a stacktail equalizer study at 12:29 PM. This study halted stacking.

MI experts began a barrier bucket study around 1:48 PM.

The TeV access ended at 2:14 PM.

The Meson access ended at 2:16 PM.

Operations put the TeV into a dry squeeze around 4 PM.

MI studies ended at 4:37 PM.

Operations prepared Pbar for an access at 4:56 PM. They were out 15 minutes later.

Operations put the TeV through a 4x0 wet squeeze around 5 PM.

Pbar began stacking at 6:13 PM.

Operations put the TeV into shot setup at 6:19 PM.

Operations established store 5276 at 8:10 PM with an initial luminosity of 152.68E30.

At 10:46 PM, a Booster kicker (MKS08) tripped off. Operations called in an expert.

The Booster expert reported that MKS08 had a bad cable.

Operations reported at midnight that the Recycler had stashed antiprotons once during the evening shift.

Tuesday March 13

The midnight shift began with Operations monitoring store 5276 (85.31E30) and the stash (45.0E10), with Pbar stacking (18.5E10), and with NuMI taking beam.

The Booster expert completed his MKS08 cable repair. Pbar resumed stacking at 12:13 AM.

At 1:13 AM, Operations reported that a Pbar Debuncher kicker (D:IKIK) was intermittently interrupting stacking.

At 6:30 AM, the Pbar tuning program was over driving a trim power supply (M:HT105). Operations disabled the program.

At 7:04 AM, power supplies for the AP2 and AP3 lines tripped off on a LCW flow indication. Operators investigated and reported a leak in a power supply (D:H717). An expert was called in to make repairs.

Operations reported at 8:35 AM that the Recycler had stashed twice during the midnight shift.

The expert completed his repair of D:H717; he replaced a broken hose. Pbar resumed stacking at 8:52 AM.

Operations established beam to the Meson MTest experiment at 9:31 AM.

At 10:41 AM, BRF3 tripped off, due to a power amplifier failure, and wouldn’t reset. Operators notified RF experts, bypassed BRF3, and brought BRF19 online to compensate.

TeV experts began an end-of-store study at 11:58 AM.

Operations terminated store 5276 at 12:16 PM.

Operations prepared the D0 collision hall for access at 12:35 PM. They came out of access at 2:50 PM.

At 3:02 PM, Operations reported that the H- Source was sparking. Operators notified an expert.

At 4:14 PM, Operations reported that the Pbar trim power supply (M:HT105) was again being overdriven by the Pbar tuning program. Operators restored the program.

At 4:19 PM, Operations reported multiple trips of HT105. Operations put a halt to stacking while experts investigated.

Preaccelerator experts reported at 4:07 PM that the H- Source was turned back on, but was having problems.

At 4:38 PM, Operations reported that M:HT102 was slowly failing; its output current was dropping and the Pbar tuning program was trying to compensate by driving the other trims harder, which was causing the M:HT105 trips.

The MTest experiment reported at 5:14 PM that they were finished with the beam for the night.

At 5:36 PM, Operations reported that the H- Source was sparking and tripped off its vacuum gauge. Operations had the source back on by 5:51 PM.

At 6:12 PM, Operations reported problems with LRF3. A Linac expert investigated.

The Linac expert completed his repair on LRF3. Beam resumed at 7:49 PM.

At 8:08 PM, a Pbar Debuncher power supply (D:H717) had no output. Operators cycled the power.

Operations put the TeV into shot setup at 8:49 PM.

At 9:38 PM, D:H717’s output went to zero for the first of three times in six minutes. Operators cycled its power each time and then called in an expert.

A power supply expert arrived at 10:22 PM to investigate the problem with D:H717.

Operations established store 5279 at 10:49 PM with an initial luminosity of 188.9E30.

At 11:13 PM, Operations reported that LRF3 was having problems. Operators changed a timing module. Beam resumed at 11:25 PM.

The Pbar power supply expert reported at 11:23 PM that he had seen the supply turn itself off. He put the supply in local and watched it for some time with no problem.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Wednesday March 14

The midnight shift began with Operations monitoring store 5279 (135.04E30) and the stash (49.6E10), with Pbar stacking (37.6E10), and with NuMI taking beam.

At 1:06 AM, LRF3 tripped off. It reset.

At 1:54 AM, the NuMI beam permit tripped off twice in 10 minutes due to power supply current drift. It reset.

At 5:04 AM, BRF7 began sparking. Operators turned off the station for 15 minutes to see if it would settle down, but it didn’t. Operators reduce the beam intensity and contacted an expert.

At 5:24 AM, BRF10 tripped off for the first of three times in 11 minutes. It reset each time.

Operations reported at 5:46 AM that with the help of an RF expert they brought BRF7 back online.

Operations established beam to MTest at 7:30 AM.

Operations reported at 8:08 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to keep store 5279 in until early evening. Due to Booster RF problems, the Run Coordinator will allow a three to four hour Booster access this afternoon. During the access, Linac will work on LRF4, and other machines will conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/14/2007    email Fermilab